OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2).

OpenStudy (bahrom7893):

okay you know the slope, but you don't know the intercept

OpenStudy (bahrom7893):

M = slope = (y-y1)/(x-x1)

OpenStudy (anonymous):

Ok. So would it be 3x-2

OpenStudy (bahrom7893):

darn it i was doing it right...

OpenStudy (bahrom7893):

nevermind, im having a long day already

OpenStudy (anonymous):

haha xD

OpenStudy (bahrom7893):

M = slope = (y-y1)/(x-x1) 3 = (-2-y1)/(1-x1)

OpenStudy (anonymous):

y=3x+1 ?

OpenStudy (bahrom7893):

okay im being retarded here. There's obviously an easy way. Final attempt at solution. If I'm wrong I'll leave this question, very embarrassed

OpenStudy (bahrom7893):

Slope intercept form: y=mx+b

OpenStudy (bahrom7893):

(1, -2); y=-2, x=1, slope = m = 3

OpenStudy (bahrom7893):

-2 = 3*1+b

OpenStudy (bahrom7893):

b = -2-3 = -5 y = 3x - 5

OpenStudy (bahrom7893):

Yea my first solution was correct.. ><

OpenStudy (anonymous):

Oh my gosh. I feel so dumb... hah. Thank you very much for your help! @bahrom7893